OpenStudy (anonymous):

Verify the trig identity: 3sin^2x + 4cos^2x = 3 + cos^2x

OpenStudy (anonymous):

\[3\sin ^{2} \theta +4\cos ^{2}\theta = 3+\cos ^{2}\theta\]

OpenStudy (anonymous):

Recall,\[4\cos ^{2}\theta =3\cos ^{2}\theta +\cos ^{2}\theta\]

OpenStudy (mathmagician):

the answer is 0

OpenStudy (anonymous):

I don't think they want to you to 'solve' the equation, just show that one side is equal to the other :)

OpenStudy (mathmagician):

Well, its the same thing- if you solve the equation and show, that the result is 0, the identity is proved

OpenStudy (anonymous):

That is true but schools have standard or conventional ways of doing it; leaving one side the same and breaking down the other.